In Children's Hospital in Seattle there are on average 60 births per week. Mothers and babies stay, on average, two days before they leave the hospital. In Swedish hospital (also in Seattle), the average number of births per week is 210. Mothers and children stay in the hospital two days on average.

a. How many new mothers are staying in Children's Hospital?

b. How many new mothers are staying in Swedish Hospital?

c. The directors of the two hospitals are negotiating unifying the children wards of the two hospitals. They believe that by doing so they will be able to reduce the number of new mothers staying in the hospital. Are they correct? How many new mothers will stay, on average, in the unified ward?
Answer #1

Hi,

Please find the detailed answer as follows:

Part A:

TR = 60 per week

TT = 2/7 weeks

WIP = TT*TR = 60*2/7 = 17.1

Total number of new mothers in the Children's Hospital = 17.1

Answer for Part A is 17.1

-------

Part B:

TR = 210 week

TT = 2/7 weeks

WIP = TT*TR = 210*2/7 = 60

Total number of new mothers in the Swedish = 60

Answer for Part B is 60

-------

Part C:

TR = 270 week

TT = 2/7 weeks

WIP = TT*TR = 270*2/7 = 77.1

Total number of new mothers in the Unified Hospital = 77.1

Answer for Part C is 77.1. No, it will not decrease the total number of mothers in the hospital.

Thanks.